# The SAGA Survey. III. A Census of 101 Satellite Systems around Milky Way-mass Galaxies

### v2

@@ Abstract
We present the third Data Release (DR3) of the Satellites Around Galactic Analogs (SAGA) Survey, a spectroscopic survey characterizing satellite galaxies around Milky Way (MW)-mass galaxies. The SAGA Survey DR3 includes 378 satellites identified across 101 MW-mass systems in the distance range 25–40.75 Mpc, and an accompanying redshift catalog of background galaxies (including about 46,000 taken by SAGA) in the SAGA footprint of 84.7 sq. deg. The number of confirmed satellites per system ranges from zero to 13, in the stellar mass range . Based on a detailed completeness model, this sample accounts for 94% of the true satellite population down to . We find that the mass of the most massive satellite in SAGA systems is the strongest predictor of satellite abundance; one-third of the SAGA systems contain LMC-mass satellites, and they tend to have more satellites than the MW. The SAGA satellite radial distribution is less concentrated than the MW, and the SAGA quenched fraction below is lower than the MW, but in both cases, the MW is within of SAGA system-to-system scatter. SAGA satellites do not exhibit a clear corotating signal as has been suggested in the MW/M31 satellite systems. Although the MW differs in many respects from the typical SAGA system, these differences can be reconciled if the MW is an older, slightly less massive host with a recently accreted LMC/SMC system.
@@ 1 Introduction
The Satellites Around Galactic Analogs (SAGA) Survey is a spectroscopic galaxy survey focusing on nearby low-mass galaxies. Its primary science goal is to characterize the satellite populations, roughly down to the luminosity of the Leo I galaxy (), around 100 Milky Way (MW) analogs outside the Local Volume (at 25–40.75 Mpc). The SAGA Survey had two previous data releases. Geha et al. (2017, Paper I hereafter) introduced the survey and showed initial results based on 27 satellites around eight MW-like systems. Mao et al. (2021, Paper II hereafter) included 127 satellites identified around 36 MW-like systems. In this work, we present the Data Release 3 (DR3) of the SAGA Survey, which includes 378 satellites identified around 101 MW-like systems. In addition, we also publish the full redshift catalog for galaxies in the SAGA fields in this data release.
@@
This paper presents the overview of the SAGA Survey, DR3 data products, and a selection of the main science results. There are two accompanying papers: Geha et al. (2024, Paper IV hereafter) will discuss in detail the star-forming properties of the SAGA satellites. Wang et al. (2024, Paper V hereafter) will present a detailed comparison with a theoretical prediction from an updated UniverseMachine galaxy–halo connection model (Behroozi et al., 2019).
This paper reviews our survey design in Section 2, including host selection, photometric catalog production, and target selection. We describe the redshift sources in Section 3, including our own observations and redshifts from the literature. We present SAGA DR3 in Section 4, which includes both the satellite sample and an overall redshift catalog. We then present the main science results in Section 5, including the stellar mass functions, quenched fraction, radial distributions, satellite abundance, and co-rotating signals. In Section 6, we discuss the MW satellite system in the context of the SAGA results, considerations for comparing SAGA results with simulations, and planned follow-up work. Readers who are familiar with previous SAGA work or want to navigate the main results of this paper quickly can read Section 7, the summary section, first.
As in Paper I and Paper II, all distance-dependent parameters are calculated assuming km s-1 Mpc-1 and . Magnitudes and colors are extinction-corrected (denoted with a subscript ‘o,’ e.g., ; using a combination of Schlegel et al. 1998 and Schlafly & Finkbeiner 2011; see Section 2.3). Absolute magnitudes are -corrected to using Chilingarian et al. (2010).
@@ 2.1 Overview of the SAGA Survey Design
The primary goal of the SAGA Survey111sagasurvey.org is to characterize the satellite galaxy populations around more than 100 Milky Way-mass galaxies down to an absolute magnitude of . As of this data release, we have completed the survey for 101 systems. To balance depth and volume, the SAGA Survey focuses on the regime slightly outside the Local Volume, selecting systems from 25 to 40.75 Mpc, as shown in Figure 1. The survey depth in apparent magnitude is roughly . The primary galaxies in these systems are selected by stellar masses and environments, which we detailed in Section 2.2. We then identify potential satellite galaxy candidates based on photometric information, using the photometric catalogs from DESI Legacy Imaging Surveys DR9 (Section 2.3, Section 2.5). We obtain redshifts for the candidates, including redshifts from existing literature, but mostly with new observations (Section 3), to confirm whether the candidates have comparable redshifts as the primary galaxies.
@@ 2.2 SAGA Host Selection and SAGA Footprint
The selection of the SAGA MW analogs (“hosts”) remains the same as described in Section 2.1 of Paper II. We briefly summarize the selection criteria here and refer the readers to Paper II for detail. Our selection is primarily based on distance, -band luminosity, and local environment. The selection criteria are:
The selection is applied on a nearby galaxy catalog that we construct from HyperLEDA (Makarov et al., 2014)222leda.univ-lyon1.fr, and referred to as the “Grand List” in Paper II. The Grand List is complete out to 60 Mpc in distance and down to (corresponding to at 60 Mpc). The -band magnitude and redshift measurements were adopted from the 2MASS Redshift Survey (2MRS; Huchra et al. 2012). The distance measurements come from a combination of the NASA–Sloan Atlas (NSA v1.0.1; Blanton et al. 2011) and redshift-independent distance measurements compiled by HyperLEDA. One of the stellar foreground selection criteria (Eq. 1c) uses the Hipparcos-2 catalog (van Leeuwen, 2007).
We also include the halo mass () estimates from the Lim et al. (2017) group catalog. Lim et al. (2017) estimated these halo masses using a group finder based on Yang et al. (2005), but with improvements to allow application on four galaxy redshift surveys. The halo mass estimation is based on both the sum of the luminosities of member galaxies and the difference in luminosity between the central and member galaxies (Lu et al., 2016), and are unbiased with a typical uncertainty of 0.2 dex when compared with hydrodynamical simulations. In the SAGA redshift range (), they estimated their group catalog is complete down to a halo mass of , which is sufficient for our use since we focus on MW-mass systems. We use this halo mass estimate in Eq. 1e and our analysis.
@@
All SAGA hosts are less massive than M 31, and most are less massive than the MW. Among the 101 hosts presented here, we identify five Local Group–like pairs (exactly two MW-mass hosts within 1 Mpc of each other), and an additional eight hosts that have precisely one MW-mass companion within 1 Mpc with the companion not being one of the 101 SAGA hosts. We indicate these hosts in Figure 3.
@@ 2.3 DESI Imaging DR9 Photometry
After SAGA Paper II, we fully updated our photometric catalog to the DESI Legacy Imaging Surveys (Dey et al., 2019) Data Release 9 (DR9), released in January 2021. All SAGA systems are within the footprint of DESI Imaging DR9. This is a major upgrade compared to our first data release, which used only SDSS DR12, and to our second data release, which used a combination of SDSS DR14, Dark Energy Survey (DES Abbott et al., 2018) DR1, and DESI Imaging DR6/7. In particular, DESI Imaging DR9 includes DECam data from other surveys, such as the DES. DESI Imaging DR9 has a single-frame median point-source detection limit in -band of (slightly different for the two sub surveys, DECaLS and BASS, Dey et al., 2019); this depth is comparable to DES and is about half magnitude deeper than SDSS.
@@
Besides shredded objects, there are also spurious objects in the catalog. The most common situations include Galactic cirri and faint stars or quasars mistaken as large galaxies when the sky background is not correctly subtracted. We remove these spurious objects with both catalog properties (mainly the flux signal-to-noise values the flags; see the function in our code base). We also remove spurious objects via visual inspection. On average, we remove about 25 spurious objects per sq. deg., which is about 0.86% of all galaxy sources down to .
In summary, our photometric catalog is based on the DESI Imaging DR9 catalog but with corrections for shredded, spurious, and misclassified objects. We ingest the photometric information including grz and WISE (W1, …, W4) bands and morphology including half-light semi-major axis () and ellipticity parameters (, ).
Because low-redshift, low-mass galaxies tend to also have lower surface brightness, we compare our photometric catalog with the SMUDGes catalog (Zaritsky et al., 2023) to ensure that low-surface-brightness galaxies (LSBGs) and ultra-diffuse galaxies (UDGs) are not systematically missing from our photometric catalog. We found 82 objects in the SMUDGes catalog that are also in the SAGA footprint. Note that the SAGA footprint is a subset of the SMUDGes footprint, but only low-surface-brightness objects are included in the SMUDGes catalog. Among the 82 SMUDGes objects, only one (0912528+351112) is not present in the SAGA photometric catalog (which is based on DESI Imaging DR9 catalog); however, this missing object has a magnitude () that is fainter than SAGA’s survey depth (). All other SMUDGes objects are present in the SAGA photometric catalog, and 14 of these are confirmed SAGA satellite galaxies (see definition in Section 4.3).
@@
2.4 GALEX photometry
@@ 2.5 Target Selection
As we demonstrated in Paper II (see the left panel of Figure 3 of Paper II), our primary targeting region selects an almost complete sample for galaxies within . The typical redshift for galaxies in the primary targeting region is around . The redshifts we have obtained since Paper II continue to confirm the effectiveness of this strategy. Nevertheless, we also continue to obtain redshifts outside the primary targeting region, especially in the region close to the boundaries of the primary targeting region. As shown in Darragh-Ford et al. (2023), low-redshift galaxies () tend to be found near these boundaries.
In summary, this target selection strategy allows us to reach a high redshift rate in the primary targeting region, where almost all galaxies reside. At the same time, we sample the rest of the photometric space, and these redshifts allow us to build a model to estimate how likely a galaxy is to be a satellite given its photometric properties. We describe this model in Section B.1, and we use it to correct for our survey incompleteness (here, by incompleteness we refer to the redshift rate less than 100%).
@@ 3 Spectroscopic Data
3.1 Multifiber Spectra: AAT and MMT
Building on our work in Paper I and Paper II, we obtained data from both the MMT/Hectospec and AAT/2dF multifiber spectrographs. In both cases, data reduction and redshift inspection are identical to that described in Paper II.
MMT/Hectospec provides 300 fibers over a 1∘ diameter field-of-view (Fabricant et al., 2005). Each fiber has a diameter of . We used Hectospec with the 270 lines mm-1 grating, resulting in wavelength coverage of 3650–9200 Å and spectral resolution of 1.2 Å pixel-1 (). Typical redshift errors are 60 km s-1. SAGA DR3 contains 19,795 unique galaxy redshifts obtained with MMT/Hectospec between 2013 and May 2022.
@@ 3.2 Single-slit Spectra: Palomar, Keck, and SALT
Single-slit spectroscopy is used to measure redshifts for galaxies that were missed in the multifiber work due to fiber collisions, low signal-to-noise, or other issues. Our primary single-slit facility is the 200-inch Hale Telescope at Palomar Observatory and the Double Spectrograph (DBSP; Oke & Gunn, 1982). We used the 316/7500 red channel grating and 600/4000 blue channel grating, resulting in wavelength coverage of 3800–9500 Å and spectral resolution of 1.1 and 1.5 Å pixel-1, respectively, for the blue and red sides. The data were reduced using PypeIt (Prochaska et al., 2020); redshifts were measured using a modified version of marz (Hinton et al., 2016). SAGA DR3 contains 458 unique redshifts obtained with Palomar/DBSP between 2019 and Sept 2022.
@@ 3.3 External Spectroscopic and HI Data Sets
Across the full SAGA footprint around 101 hosts, SDSS contributed about 13,000 galaxy redshifts. DESI, GAMA, PRIMUS, and NSA each contributes around 3,500–5,000 galaxy redshifts. The remaining surveys listed above each contribute hundreds to 1,500 galaxy redshifts, and around 6,000 in total. We only ingested literature redshifts with high confidence; however, in some rare cases, a literature redshift can be inaccurate. When a literature redshift indicates a satellite galaxy (about 40% of the SAGA satellites have literature redshifts), we retarget the galaxy to verify the literature velocity measurement and to ensure homogeneity, but we do not count this case as a SAGA discovery.
We also ingested Hi data for galaxies that are covered by the Arecibo Legacy Fast ALFA Survey (ALFALFA; Haynes et al., 2018) and the FAST all sky HI Survey (FASHI; Zhang et al., 2024). Due to the large beam size (3′) of Hi measurement, we use the photometric counterpart information available in these catalog to assist the match. When the photometric counterpart information is not available or has a low confidence, we adopt the Hi measurements if a nearby source in the SAGA catalog has a consistent redshift. We also matched our catalog to the HI Parkes All-Sky Survey (HIPASS; Meyer et al., 2004); however, due to an even larger beam size (15′), we only search around confirmed satellites and included HIPASS data for two of the SAGA satellites.

### v3

@@ Abstract
We present Data Release 3 (DR3) of the Satellites Around Galactic Analogs (SAGA) Survey, a spectroscopic survey characterizing satellite galaxies around Milky Way (MW)-mass galaxies. The SAGA Survey DR3 includes 378 satellites identified across 101 MW-mass systems in the distance range of 25–40.75 Mpc, and an accompanying redshift catalog of background galaxies (including about 46,000 taken by SAGA) in the SAGA footprint of 84.7 sq. deg. The number of confirmed satellites per system ranges from zero to 13, in the stellar mass range of . Based on a detailed completeness model, this sample accounts for 94% of the true satellite population down to . We find that the mass of the most massive satellite in SAGA systems is the strongest predictor of satellite abundance; one-third of the SAGA systems contain LMC-mass satellites, and they tend to have more satellites than the MW. The SAGA satellite radial distribution is less concentrated than the MW’s, and the SAGA quenched fraction below is lower than the MW’s, but in both cases, the MW is within of SAGA system-to-system scatter. SAGA satellites do not exhibit a clear corotating signal as has been suggested in the MW/M31 satellite systems. Although the MW differs in many respects from the typical SAGA system, these differences can be reconciled if the MW is an older, slightly less massive host with a recently accreted LMC/SMC system.
@@ 1 Introduction
The Satellites Around Galactic Analogs (SAGA) Survey is a spectroscopic galaxy survey focusing on nearby low-mass galaxies. Its primary science goal is to characterize the satellite populations, roughly down to the luminosity of the Leo I galaxy (), around 100 Milky Way (MW) analogs outside the Local Volume (at 25–40.75 Mpc). The SAGA Survey had two previous data releases. Geha et al. (2017, hereafter Paper I) introduced the survey and showed initial results based on 27 satellites around eight MW-like systems. Mao et al. (2021, hereafter Paper II) included 127 satellites identified around 36 MW-like systems. In this work, we present Data Release 3 (DR3) of the SAGA Survey, which includes 378 satellites identified around 101 MW-like systems. In addition, we also publish the full redshift catalog for galaxies in the SAGA fields in this data release.
@@
This paper presents the overview of the SAGA Survey, DR3 data products, and a selection of the main science results. There are two accompanying papers: Geha et al. (2024, hereafter Paper IV) will discuss in detail the star-forming properties of the SAGA satellites. Wang et al. (2024, hereafter Paper V) will present a detailed comparison with a theoretical prediction from an updated UniverseMachine galaxy–halo connection model (Behroozi et al., 2019).
This paper reviews our survey design in Section 2, including host selection, photometric catalog production, and target selection. We describe the redshift sources in Section 3, including our own observations and redshifts from the literature. We present SAGA DR3 in Section 4, which includes both the satellite sample and an overall redshift catalog. We then present the main science results in Section 5, including the stellar mass functions (SMFs), quenched fraction, radial distributions, satellite abundance, and co-rotating signals. In Section 6, we discuss the MW satellite system in the context of the SAGA results, considerations for comparing SAGA results with simulations, and planned follow-up work. Readers who are familiar with previous SAGA work or want to navigate the main results of this paper quickly can read Section 7, the summary section, first.
As in Paper I and Paper II, all distance-dependent parameters are calculated assuming km s-1 Mpc-1 and . Magnitudes and colors are extinction-corrected (denoted with a subscript ‘o,’ e.g., ; using a combination of data from Schlegel et al. 1998 and Schlafly & Finkbeiner 2011; see Section 2.3). Absolute magnitudes are -corrected to using the formulas from Chilingarian et al. (2010).
@@ 2.1 Overview of the SAGA Survey Design
The primary goal of the SAGA Survey111sagasurvey.org is to characterize the satellite galaxy populations around more than 100 Milky Way-mass galaxies down to an absolute magnitude of . As of this data release, we have completed the survey for 101 systems. To balance depth and volume, the SAGA Survey focuses on the regime slightly outside the Local Volume, selecting systems from 25 to 40.75 Mpc, as shown in Figure 1. The survey depth in apparent magnitude is roughly . The primary galaxies in these systems are selected by stellar masses and environments, which we detail in Section 2.2. We then identify potential satellite galaxy candidates based on photometric information, using the photometric catalogs from DESI Legacy Imaging Surveys DR9 (Section 2.3, Section 2.5). We obtain redshifts for the candidates, including redshifts from existing literature, but mostly with new observations (Section 3), to confirm whether the candidates have comparable redshifts as the primary galaxies.
@@ 2.2 SAGA Host Selection and SAGA Footprint
The selection of the SAGA MW analogs (“hosts”) remains the same as described in Section 2.1 of Paper II. We briefly summarize the selection criteria here and refer the readers to Paper II for details. Our selection is primarily based on distance, -band luminosity, and local environment. The selection criteria are
The selection is applied on a nearby galaxy catalog that we constructed from HyperLEDA (Makarov et al., 2014)222leda.univ-lyon1.fr, and referred to as the “Grand List” in Paper II. The Grand List is complete out to 60 Mpc in distance and down to (corresponding to at 60 Mpc). The -band magnitude and redshift measurements were adopted from the 2MASS Redshift Survey (2MRS; Huchra et al. 2012). The distance measurements come from a combination of the NASA–Sloan Atlas (NSA v1.0.1; Blanton et al. 2011) and redshift-independent distance measurements compiled by HyperLEDA. One of the stellar foreground selection criteria (Eq. 1c) uses the Hipparcos-2 catalog (van Leeuwen, 2007).
We also include the halo mass () estimates from the Lim et al. (2017) group catalog. Lim et al. (2017) estimated these halo masses using a group finder based on the work of Yang et al. (2005), but with improvements to allow application on four galaxy redshift surveys. The halo mass estimation is based on both the sum of the luminosities of member galaxies and the difference in luminosity between the central and member galaxies (Lu et al., 2016), and is unbiased with a typical uncertainty of 0.2 dex when compared with hydrodynamical simulations. In the SAGA redshift range (), they estimated their group catalog is complete down to a halo mass of , which is sufficient for our use since we focus on MW-mass systems. We use this halo mass estimate in Eq. 1e and our analysis.
@@
All SAGA hosts are less massive than M 31, and most are less massive than the MW. Among the 101 hosts presented here, we identify five “Local Group–like” pairs (exactly two MW-mass hosts within 1 Mpc of each other), and an additional eight hosts that have precisely one MW-mass companion within 1 Mpc with the companion not being one of the 101 SAGA hosts. We indicate these hosts in Figure 3.
@@ 2.3 DESI Imaging DR9 Photometry
After SAGA Paper II, we fully updated our photometric catalog to the DESI Legacy Imaging Surveys (Dey et al., 2019) Data Release 9 (DR9), released in January 2021. All SAGA systems are within the footprint of DESI Imaging DR9. This is a major upgrade compared to our first data release, which used only SDSS DR12, and to our second data release, which used a combination of SDSS DR14, Dark Energy Survey (DES Abbott et al., 2018) DR1, and DESI Imaging DR6/7. In particular, DESI Imaging DR9 includes DECam data from other surveys, such as the DES. DESI Imaging DR9 has a single-frame median point-source detection limit in the band of (slightly different for the two sub-surveys, DECaLS and BASS, Dey et al., 2019); this depth is comparable to DES and is about half magnitude deeper than SDSS.
@@
Besides shredded objects, there are also spurious objects in the catalog. The most common situations include Galactic cirri and faint stars or quasars mistaken as large galaxies when the sky background is not correctly subtracted. We remove these spurious objects with both catalog properties (mainly the flux signal-to-noise values and the flags; see the function in our code base). We also remove spurious objects via visual inspection. On average, we remove about 25 spurious objects per sq. deg., which is about 0.86% of all galaxy sources down to .
In summary, our photometric catalog is based on the DESI Imaging DR9 catalog but with corrections for shredded, spurious, and misclassified objects. We ingest the photometric information, including grz and WISE (W1, …, W4) bands, and morphology, including half-light semi-major axis () and ellipticity parameters (, ).
Because low-redshift, low-mass galaxies tend to also have lower surface brightness, we compare our photometric catalog with the SMUDGes catalog (Zaritsky et al., 2023) to ensure that low-surface-brightness galaxies (LSBGs) and ultra-diffuse galaxies (UDGs) are not systematically missing from our photometric catalog. We found 82 objects in the SMUDGes catalog that are also in the SAGA footprint. Note that the SAGA footprint is a subset of the SMUDGes footprint, but only low-surface-brightness objects are included in the SMUDGes catalog. Among the 82 SMUDGes objects, only one (0912528+351112) is not present in the SAGA photometric catalog (which is based on the DESI Imaging DR9 catalog); however, this missing object has a magnitude () that is fainter than SAGA’s survey depth (). All other SMUDGes objects are present in the SAGA photometric catalog, and 14 of these are confirmed SAGA satellite galaxies (see definition in Section 4.3).
@@
2.4 Galaxy Evolution Explorer photometry
@@ 2.5 Target Selection
As we demonstrated in Paper II (see the left panel of Figure 3 in Paper II), our primary targeting region selects an almost complete sample for galaxies within . The typical redshift for galaxies in the primary targeting region is around . The redshifts we have obtained since Paper II continue to confirm the effectiveness of this strategy. Nevertheless, we also continue to obtain redshifts outside the primary targeting region, especially in the region close to the boundaries of the primary targeting region. As shown in Darragh-Ford et al. (2023), low-redshift galaxies () tend to be found near these boundaries.
In summary, this target selection strategy allows us to reach a high redshift rate in the primary targeting region, where almost all galaxies reside. At the same time, we sample the rest of the photometric space, and these redshifts allow us to build a model to estimate how likely a galaxy is to be a satellite, given its photometric properties. We describe this model in Section B.1, and we use it to correct for our survey incompleteness (here, by incompleteness we refer to the redshift rate less than 100%).
@@ 3 Spectroscopic Data
3.1 Multifiber Spectra: Anglo-Australian Telescope and MMT
Building on our work in Paper I and Paper II, we obtained data from both the MMT/Hectospec and Anglo-Australian Telescope (AAT)/2dF multifiber spectrographs. In both cases, data reduction and redshift inspection are identical to that described in Paper II.
MMT/Hectospec provides 300 fibers over a 1∘ diameter field of view (Fabricant et al., 2005). Each fiber has a diameter of . We used Hectospec with the 270 line mm-1 grating, resulting in wavelength coverage of 3650–9200 Å and spectral resolution of 1.2 Å pixel-1 (). Typical redshift errors are 60 km s-1. SAGA DR3 contains 19,795 unique galaxy redshifts obtained with MMT/Hectospec between 2013 and May 2022.
@@ 3.2 Single-slit Spectra: Palomar, Keck, and SALT
Single-slit spectroscopy is used to measure redshifts for galaxies that were missed in the multifiber work due to fiber collisions, low signal-to-noise, or other issues. Our primary single-slit facility is the 200-inch Hale Telescope at the Palomar Observatory and the Double Spectrograph (DBSP; Oke & Gunn, 1982). We used the 316/7500 red channel grating and 600/4000 blue channel grating, resulting in wavelength coverage of 3800–9500 Å and spectral resolution of 1.1 and 1.5 Å pixel-1, respectively, for the blue and red sides. The data were reduced using PypeIt (Prochaska et al., 2020); redshifts were measured using a modified version of Marz (Hinton et al., 2016). SAGA DR3 contains 458 unique redshifts obtained with Palomar/DBSP between 2019 and Sept 2022.
@@ 3.3 External Spectroscopic and HI Data Sets
Across the full SAGA footprint around 101 hosts, SDSS contributed about 13,000 galaxy redshifts. DESI, GAMA, PRIMUS, and NSA each contribute around 3,500–5,000 galaxy redshifts. The remaining surveys listed above each contribute hundreds to 1,500 galaxy redshifts, and around 6,000 in total. We only ingested literature redshifts with high confidence; however, in some rare cases, a literature redshift can be inaccurate. When a literature redshift indicates a satellite galaxy (about 40% of the SAGA satellites have literature redshifts), we retarget the galaxy to verify the literature velocity measurement and to ensure homogeneity, but we do not count this case as a SAGA discovery.
We also ingested Hi data for galaxies that are covered by the Arecibo Legacy Fast ALFA Survey (ALFALFA; Haynes et al., 2018) and the FAST all sky HI Survey (FASHI; Zhang et al., 2024). Due to the large beam size (3′) of Hi measurement, we use the photometric counterpart information available in these catalogs to assist the match. When the photometric counterpart information is not available or has a low confidence, we adopt the Hi measurements if a nearby source in the SAGA catalog has a consistent redshift. We also matched our catalog to the HI Parkes All-Sky Survey (HIPASS; Meyer et al., 2004); however, due to an even larger beam size (15′), we only search around confirmed satellites and included HIPASS data for two of the SAGA satellites.